OpenStudy (anonymous):

What is the vertex of the parabola given by y = x2 − 10x + 17?PLEASE HELP ME!!!!!!!!!

OpenStudy (anonymous):

Parabola is in standard form. Put into vertex form. y(x)+25=x^2-10x+25+17 y(x)+25=x^2-10x+25+17 y(x)+25= (x-5)^2 + 17 y(x)= (x-5)^2 -8 Vertex form formula is y= a(x-h)2+k -h,k is coordinate when in vertex form. (5,-8)